# Compress then Serve: Serving Thousands of LoRA Adapters with Little Overhead

### v1

@@ Abstract
Fine-tuning large language models (LLMs) with low-rank adapters (LoRAs) has become common practice, often yielding numerous copies of the same LLM differing only in their LoRA updates. This paradigm presents challenges for systems that serve real-time responses to queries that each involve a different LoRA. Prior works optimize the design of such systems but still require continuous loading and offloading of LoRAs, as it is infeasible to store thousands of LoRAs in GPU memory. To mitigate this issue, we investigate the efficacy of compression when serving LoRA adapters. We consider compressing adapters individually via SVD and propose a method for joint compression of LoRAs into a shared basis paired with LoRA-specific scaling matrices. Our experiments with up to 500 LoRAs demonstrate that compressed LoRAs preserve performance while offering major throughput gains in realistic serving scenarios with over a thousand LoRAs, maintaining 75% of the throughput of serving a single LoRA.
@@ 1 Introduction
The myriad uses for foundation models (FMs) has led to a proliferation of specialized models, each fine-tuned to perform downstream tasks. The growing number of parameters of these models, however, incurs significant costs both for the fine-tuning and for serving these models.
To avoid fine-tuning giant foundation models, parameter-efficient fine-tuning (PEFT) algorithms update a smaller set of values that define an edit to the original model. An especially successful PEFT method is low-rank adaptation (LoRA) [8], which achieves parameter-efficient fine-tuning by learning low-rank additive changes to neural network matrices. Because of the low-rank parameterization, these matrices (called adapter weights) contain orders-of-magnitude fewer parameters than the base model. LoRA can achieve performance on par with full fine-tuning [8]. In addition, by merging the adapter weights into the base model, the fine-tuned model does not suffer from increased latency.
LoRA’s popularity has triggered a growing need to serve large collections of LoRA adapters at the scale where serving multiple copies of the base model becomes infeasible [25]. In the extreme, one may wish to serve hundreds to thousands of LoRAs, where each incoming request requires a different LoRA adapter, e.g., for a personalized chat assistant. To this end, S-LoRA [25] separates the base model computation from individual LoRA computations and optimizes the system’s inner workings via custom CUDA kernels and memory management to increase the throughput when serving multiple LoRAs. Despite the optimized system design, serving LoRAs still has a fundamental limitation: when the number of adapters is large, they need to be constantly loaded and offloaded from GPU memory to accommodate incoming requests, degrading throughput.
In this work, we consider the problem of compressing a collection of LoRAs. We have two key objectives: (1) preserving the performance of the original LoRAs and (2) improving the throughput of serving many LoRAs. We formulate LoRA compression as a reconstruction problem, where the goal is to approximate the original adapters via collections of matrices of a smaller total size. We investigate two approaches: compressing each LoRA individually by lowering its rank via SVD and compressing LoRAs jointly by finding a shared basis and LoRA-specific scaling matrices. The former is inspired by [11], who demonstrated that lowering LoRA ranks is beneficial for multi-task learning and model merging. For the latter, we propose a joint compression algorithm inspired by models for joint diagonalization.
@@
Our individual and joint compression methods present interesting trade-offs. For example, joint compression can achieve extreme parameter count reduction when representing any specific LoRA by using a shared basis that can be pre-loaded onto the GPU. The shared basis, however, will typically increase the rank of LoRAs. On the other hand, individual compression can reduce the rank, but is limited in terms of its compression efficiency. We investigate how these trade-offs affect throughput when serving 1000s of LoRAs using vLLM [12], a state-of-the-art LLM serving system. In Figure 1 we illustrate the benefits of compression and the effect of these trade-offs. When serving over 1000 LoRAs, compression increases throughput over 2 times and maintains 75% of the throughput of serving the base LLM (or a single LoRA merged into the LLM).
@@
  • We formulate the problem of compressing a collection of LoRAs and propose two solutions: individual and joint compression.
@@
  • We train a collection of 500 high-quality LoRAs for Mistral-7B-Instruct-v0.2 [10] on 500 natural instruction tasks [27] and demonstrate that our compression techniques can preserve the performance of the original LoRAs. We will release the 500 LoRAs to facilitate future work on LoRA compression as well as the code to compress new models.
@@
  • We analyze the trade-offs between individual and joint compression, presenting opportunities for further improvement of LLM serving systems.
@@ 2 Related Work
Parameter-efficient fine-tuning (PEFT) has become a prevalent method for updating foundation models thanks to the need for efficiency in training and communication [14]. Many PEFT methods have been proposed, e.g. [7, 16], but the most used in practice is LoRA [8], partially due to the ease of switching between LoRAs in inference time.
Several works propose improvements to LoRA [17, 26], sometimes with algebraic methods like SVD [19, 34, 11] or by leveraging its statistical properties [35, 33]. Relatively few methods, however, accelerate inference times. S-LoRA [25] provides an efficient means of switching between LoRAs. F-LoRA [28] adapts training to reduce batch multiplications and thus to accelerate inference. Our method achieves a similar outcome (see Appendix E) without changing the LoRA formulation or requiring that LoRAs be trained in a dedicated way; from this perspective, an advantage of our work is that future generic improvements to LoRA will also benefit (e.g., [19]).
There are many efforts to compress models [1, 5, 24, 13]—including some that consider LoRAs specifically—to accelerate inference. Predominantly, pruning and sparsification methods delete some of the weights [31], and quantization methods reduce the precision of the weights [4]. Some works compress weights to reduce model size but typically require decompression and hence do not save GPU memory [6]. Similarly to our work, while most methods increase speed at the cost of performance, a few note increased performance and generalization after compression [31, 21, 6, 24].
Our work also relates to model merging [3, 30, 18] and mixture of experts methods [20]. Such methods reuse models trained by others [2, 22], serving them together as one compressed model. Despite this similarity, these methods create a single general model that acts on any input, while our model allows for more performant per-task solutions.
@@ 3 Rank-Based LoRA Compression
In formulating our compression algorithms, we consider a collection of given LoRA adapters $\{(A_{i},B_{i})\}_{i=1}^{n}$ that we would like to serve. We let $r_{i}$ refer to the rank of the LoRA adapter-pair $(A_{i},B_{i})$, i.e., $B_{i}\in\mathbb{R}^{d_{B}\times r_{i}}$, $A_{i}\in\mathbb{R}^{r_{i}\times d_{A}}$. After laying out the basic considerations in designing a compression scheme in §3.1, we consider a simple baseline parameter reduction approach §3.2 as well as a more aggressive joint diagonalization method in §3.3.
3.1 Desiderata
@@
3.2 Independent Compression via SVD
The simplest approach is to replace each rank-$r_{i}$ LoRA adapter $B_{i}A_{i}$ with a reduced rank-$r$ approximation, where typically $r<\frac{1}{n}\sum_{i=1}^{n}r_{i}$:
$\operatorname{SVD}_{r}(B_{i}A_{i})=U_{i}\Sigma_{i}V_{i}^{\top},\quad\forall i=1,\dots,n~{}~{}~{}~{}~{}~{}(\mathsf{r-SVD})$ (1)
Thanks to the Eckart-Young Theorem, this strategy finds the rank-$r$ approximation that best fits $B_{i}A_{i}$ in terms of Frobenius norm. As $\Sigma_{i}V_{i}^{\top}$ can be saved as a single matrix, this approach reduces the number of parameters used from $\sum_{i}(d_{A}+d_{B})r_{i}$ to $rn(d_{A}+d_{B})$. While this approximation is effective, it is limited in its compression abilities as $n$ increases, since it does not share information across LoRAs. This shortcoming motivates our next approach.
@@ 3.3 Joint Diagonalization
Next, we suggest a Joint Diagonalization (JD) method, which optimizes a shared basis onto which we can project the set of $n$ LoRAs. This will allow structure to be shared, implicitly clustering the collection of LoRAs.
In this model, each LoRA-product $B_{i}A_{i}$ is factorized into the form $U\Sigma_{i}V$, where $U$ and $V$ are shared across all LoRAs and $\Sigma_{i}$ is specific to each LoRA. In this formulation, every $\Sigma_{i}$ shares the same rank $r$. This allows $U$ and $V$ to be pre-loaded onto the GPU, with $\Sigma_{i}$ loaded when necessary for each batch. The matrices $\Sigma_{i}$ can be either diagonal or small square matrices, accelerating the forward pass compared to conventional multi-LoRA serving configurations.
@@
Using the Frobenius norm has the added benefit of making the objective convex in each argument separately, suggesting the possibility of efficient optimization. This objective function is underdetermined, however, so we consider two constrained regimes below.
Full $\Sigma_{i}$ approximation. The first method we call $\mathsf{JD-Full}$. Without loss of generality, $U$ and $V$ can be constrained to be orthogonal, so long as $\Sigma_{i}$ remains an unconstrained full matrix. $\mathsf{JD-Full}$ adopts this restriction to make the optimization better posed, but note it does not restrict the expressiveness of the objective (2). This setting yields the following optimization problem:
@@
An efficient alternating algorithm to solve this objective function can be found in Appendix B.
@@
An efficient alternating least squares algorithm to optimize this objective can be found in Appendix B. This diagonal version has some per-LoRA parameter savings when compared to $\mathsf{JD-Full}$, since the diagonal $\Sigma_{i}$ only needs $r$ parameters instead of $r^{2}$.
@@ 4 Theoretical Analysis
While SVD-based approaches are relatively well-understood, in this section, we seek to better understand the role of the joint diagonalization method presented in §3.3. We will focus on the full-$\Sigma_{i}$ case with orthogonal $U$, $V$ matrices. Note that, for the same $r$, the $r\!\!-\!\!\mathsf{JD\!-\!Diag}$ has at least as large reconstruction error as $r\!\!-\!\!\mathsf{JD\!-\!Full}$ since it imposes an additional constraint on the $\Sigma_{i}$.
Firstly, note that perfect reconstruction can be achieved if and only if $r$ is large enough:
@@
Note $\max_{i}r_{i}\leq\tilde{r}\leq\sum_{i=1}^{n}r_{i}$. Then $\mathsf{JD-Full}$ (3) with $r=\tilde{r}$ achieves lossless compression (perfect reconstruction), and using $r<\tilde{r}$ will give nonzero reconstruction error.
Proof.
There exist $U,V$ such that all the $B_{i}$, $A_{i}$ are in the spans of $U,V$ resp. if and only if $r\geq\tilde{r}$. ∎
@@
This reconstruction loss can be significant, as the following theorem shows (proved in Appendix C):
@@
Let $\sigma_{j}$ be the singular values of $L$, sorted from largest to smallest, and let $\bar{\sigma}_{j}$ be the singular values of $\sum_{i=1}^{n}B_{i}A_{i}$. Then, using $\mathsf{JD-Full}$ (3),
@@
Remark 2 (Upper bound and clustering).
@@
Suppose (e.g., due to normalization) that the inputs to the joint diagonalization algorithm all have unit Frobenius norm, i.e., $\|B_{i}A_{i}\|_{\mathrm{Fro}}=1$. Moreover, assume that the LoRAs are all orthogonal in the sense $\mathrm{tr}((B_{i}A_{i})(B_{j}A_{j})^{\top})=0$ for $i\neq j$. Then, using the $\mathsf{JD-Full}$ method (3), we have $1\leq\sum_{i=1}^{n}\|\Sigma_{i}\|^{2}_{\mathrm{Fro}}\leq\min(r^{2},n),$ implying that the sum of squared Frobenius norms of the reconstructed LoRAs satisfies
@@
Note that since the loss space of transformers is highly complex, increasing reconstruction error for the weights does not necessarily imply degrading ability to approximate the desired LLM outputs. Interestingly, in an experiment shown in Figure 4 below, we indeed see that while large reconstruction error rapidly decreases performance, moderate (but still relatively large, at around 0.4) reconstruction error does not damage performance and may even slightly outperform the zero reconstruction error setting. This observation motivates our focus on minimizing weight reconstruction error, while also pointing towards the possibility that our approach is capable of achieving something deeper than simply low-loss compression. Specifically, the tendency of the Frobenius-norm based joint diagonalization is to find subspaces that are shared with many LoRAs when $r$ is large, and merge subspaces together towards the mean when $r$ is small. When $r$ is severely below the natural rank of the problem, this tendency towards averaging all or some of the LoRA signals directly connects to the concept of merging LoRAs, whose empirical success in works such as [23, 9] could explain the success of our procedure despite the nonlinearity of the underlying transformers.
Experiments in Appendix F.10 explore this idea further, comparing reconstruction of real-world LoRAs to reconstruction of randomly sampled LoRAs. The reconstruction error is generally large, but significantly lower than the reconstruction error for random noise, indicating that there is a major shared component between the LoRAs that is being successfully retained.

### v2

@@ Abstract
Fine-tuning large language models (LLMs) with low-rank adaptations (LoRAs) has become common practice, often yielding numerous copies of the same LLM differing only in their LoRA updates. This paradigm presents challenges for systems that serve real-time responses to queries that each involve a different LoRA. Prior works optimize the design of such systems but still require continuous loading and offloading of LoRAs, as it is infeasible to store thousands of LoRAs in GPU memory. To mitigate this issue, we investigate the efficacy of model compression when serving LoRAs. We propose a method for joint compression of LoRAs into a shared basis paired with LoRA-specific scaling matrices. We extend our algorithm to learn clusters of LoRAs that are more amenable to joint compression, allowing it to scale gracefully to large LoRA collections. Our experiments with up to 500 LoRAs demonstrate that compressed LoRAs preserve performance while offering major throughput gains in realistic serving scenarios with over a thousand LoRAs, maintaining 80% of the throughput of serving a single LoRA.
@@ 1 Introduction
The myriad uses for foundation models (FMs) have led to a proliferation of specialized models, each fine-tuned to perform a downstream task. To avoid fine-tuning foundation models with billions of parameters, more parameter-efficient fine-tuning (PEFT) algorithms were proposed. An especially successful PEFT method is low-rank adaptation (LoRA) (Hu et al., 2021), which learns low-rank additive changes to neural network matrices. Because of the low-rank parameterization, these matrices (called adapter weights) contain orders-of-magnitude fewer parameters than the base model. Still, LoRA can achieve performance on par with full fine-tuning (Hu et al., 2021).
LoRA’s popularity has triggered a growing need to serve large collections of LoRA adapters at scale. For example, proprietary and open-source LLM providers offer fine-tuning services (OpenAI, 2024; TogetherAI, 2024; Predibase, 2024) with user bases likely in thousands or even hundreds of thousands. As each user wants to use their own fine-tuned version of the LLM, simply serving a dedicated fine-tuned LLM per user becomes infeasible. To this end, S-LoRA (Sheng et al., 2023) considers a system where only the base LLM is placed on an inference server and individual LoRA adapters are switched as needed at inference time. S-LoRA optimizes the system’s inner workings via custom CUDA kernels and memory management to increase the throughput when serving multiple LoRAs. Such multi-LoRA system design has also been adopted in vLLM (Kwon et al., 2023), a state-of-the-art LLM serving engine. Despite the optimized system design, serving LoRAs still has a fundamental limitation: when the number of adapters is large, they need to be constantly loaded and offloaded from GPU memory to accommodate incoming requests, degrading throughput.
The problem of accommodating multiple LoRA adapters is also apparent when placing LLMs on edge devices, where smaller LLMs are fine-tuned for various tasks and the adapters are swapped depending on the task at hand (Gunter et al., 2024). In this case, the amount of adapters is smaller, e.g., a few dozen (Gunter et al., 2024), but the memory constraints are also tighter due to the limited capacity of edge devices.
In this work, we consider the problem of compressing a collection of LoRAs. We have two key objectives: (1) preserving the performance of the original LoRAs and (2) improving the throughput of serving many LoRAs. We formulate LoRA compression as a reconstruction problem, where the goal is to approximate the original adapters via collections of matrices of a smaller total size. We investigate an approach based on compressing LoRAs jointly by finding a shared basis and LoRA-specific scaling matrices, and propose a joint diagonalization-based algorithm (JD). To improve reconstruction error for large numbers of LoRAs while keeping the number of parameters in check, we propose a clustering approach where each cluster is compressed independently using the joint diagonalization algorithm. Our clustering algorithm is based on alternating between optimizing the cluster assignments and the per-cluster reconstruction error.
@@
We showcase the benefits of joint compression in Figure 1. We chose JD configurations that match the constraints. When serving up to 64 unique LoRAs we use JD without clustering and for 128 or more, we pick the number of clusters to match the performance of compressed and original LoRAs. In each case, the GPU memory footprint of the compressed and original LoRAs is matched for a fair comparison to the vLLM’s multi-LoRA inference engine. When serving over 1000 LoRAs, compression increases throughput by a factor of 1.6 and maintains 80% of the throughput of serving the base LLM (or a single LoRA merged into the LLM). Detailed results are presented in Section 6.
@@
  • We formulate the problem of compressing a collection of LoRAs and propose a joint compression scheme based on joint diagonalization.
  • For large numbers of LoRAs, we scale the joint compression scheme by proposing a clustering algorithm where each cluster is jointly compressed to minimize the overall reconstruction error.
@@
  • We train a collection of 500 high-quality LoRAs for Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a) on 500 natural instruction tasks (Wang et al., 2022) and demonstrate that our compression techniques preserve the performance of the original LoRAs. We will release the 500 LoRAs to facilitate future work on LoRA compression as well as the code for our method.
@@ 2 Related Work
Parameter-efficient fine-tuning (PEFT) has become prevalent for updating foundation models thanks to the need for efficiency in training and communication (Lialin et al., 2023). Many PEFT methods have been proposed, e.g. (Houlsby et al., 2019; Liu et al., 2022) and LoRA (Hu et al., 2021) became the standard, partially due to the ease of switching between LoRAs in inference time.
Several works improve LoRA (Liu et al., 2024; Wang et al., 2024), sometimes with algebraic methods like SVD (Meng et al., 2024; Zhang et al., 2023; Jiang et al., 2023b) or by leveraging its statistical properties (Zhu et al., 2024; Zeng & Lee, 2024). Relatively few, however, accelerate inference times. S-LoRA (Sheng et al., 2023) provides an efficient means of switching between LoRAs. F-LoRA (Wen & Chaudhuri, 2024) adapts training to reduce batch multiplications, accelerating inference. Our method achieves a similar outcome (see Appendix D) without changing the LoRA formulation or requiring that LoRAs be trained in a dedicated way; future improvements to LoRA will also benefit from this aspect of our work (e.g., Meng et al. (2024)).
There are many efforts to compress models (Cheng et al., 2017; Gholami et al., 2022; Sharma et al., 2024; Li et al., 2018)—including some specifically for LoRAs—to accelerate inference. Predominantly, pruning and sparsification methods delete weights (Yadav et al., 2023a), and quantization methods reduce the weights’ precision (Dettmers et al., 2024). Some works compress weights to reduce model size but typically require decompression and hence do not save GPU memory (Hershcovitch et al., 2024). Similarly to our work, while most methods increase speed at the cost of performance, a few note increased performance and generalization after compression (Yadav et al., 2023a; Nadjahi et al., 2023; Hershcovitch et al., 2024; Sharma et al., 2024).
Our work also relates to model merging (Choshen et al., 2022; Wortsman et al., 2022; Matena & Raffel, 2021) and mixture of experts methods (Muqeeth et al., 2024; Yadav et al., 2024). These methods reuse models trained by others (Choshen et al., 2023; Raffel, 2023), serving them together as one compressed model. Despite this similarity, these methods create a single general model that acts on any input, while our model allows for more performant per-task solutions.
@@ 3 Rank-Based LoRA Compression
In formulating our compression algorithms, we consider a collection of given LoRA adapters $\{(A_{i},B_{i})\}_{i=1}^{n}$ that we would like to serve. We let $r_{i}$ refer to the rank of the LoRA adapter-pair $(A_{i},B_{i})$, i.e., $B_{i}\in\mathbb{R}^{d_{B}\times r_{i}}$, $A_{i}\in\mathbb{R}^{r_{i}\times d_{A}}$.
@@ 3.1 Joint Diagonalization
For compression to scale to large numbers of LoRAs, the compressed number of parameters should not scale linearly with $n$. Hence seeking to compress each LoRA individually (e.g., via SVD as detailed in the experimental baselines) is inherently limited.
To address this, we suggest a Joint Diagonalization (JD) method, which optimizes a shared basis onto which we can project the set of $n$ LoRAs. This will allow structure to be shared, implicitly grouping and/or merging the collection of LoRAs.
In this model, each LoRA product $B_{i}A_{i}$ is factorized into the form $U\Sigma_{i}V$, where $U$ and $V$ are shared across all LoRAs and $\Sigma_{i}$ is specific to each LoRA. In this formulation, every $\Sigma_{i}$ shares the same rank $r$. This allows $U$ and $V$ to be pre-loaded onto the GPU, with $\Sigma_{i}$ loaded when necessary for each batch. The matrices $\Sigma_{i}$ can be either diagonal or small square matrices, accelerating the forward pass compared to conventional multi-LoRA serving configurations.
@@
Note this problem is not solved by a single matrix SVD, since $U$ and $V$ are shared among all terms but the $\Sigma_{i}$’s are not. Using the Frobenius norm has the added benefit of making the objective convex in each argument separately, suggesting the possibility of efficient optimization. This objective function is underdetermined, however, so we consider two constrained regimes below.
Full $\Sigma_{i}$ approximation. The first method we call $\mathsf{JD-Full}$. Without loss of generality, $U$ and $V$ can be constrained to be orthogonal, so long as $\Sigma_{i}$ remains an unconstrained full matrix. $\mathsf{JD-Full}$ adopts this restriction to make the optimization better posed, but note it does not restrict the expressiveness of the objective equation 1. This setting yields the following optimization problem:
@@
An efficient alternating algorithm to solve this objective function can be found in Appendix A.
@@
An efficient alternating least squares algorithm to optimize this objective can be found in Appendix A. This diagonal version has some per-LoRA parameter savings when compared to $\mathsf{JD-Full}$, since the diagonal $\Sigma_{i}$ only needs $r$ parameters instead of $r^{2}$.
3.2 Clustering
As the number of LoRAs $n$ grows and becomes more diverse, the rank $r$ needed for Joint Diagonalization to achieve good performance will tend to increase. This increases the size and number of parameters of each $\Sigma_{i}$ that needs to be stored, especially for $\operatorname{JD-Full}$ which will require $O(nr^{2})$ storage for these matrices. If the necessary $r$ is growing proportionally to $n$, then this storage will eventually become the bottleneck.
To resolve this limitation with very large $n$, we propose to group the $n$ LoRAs into $|C_{j}|$ clusters $C_{j}$. Each cluster is given its own rank $r$ JD compression, and the clusters are chosen such that the overall reconstruction error is minimized. Specifically, the overall objective is
$\min_{\{\{C_{j}\},{U_{j},V_{j}}\},\{\Sigma_{i}\}}\sum_{j}\sum_{i\in C_{j}}||B_{i}A_{i}-U_{j}\Sigma_{i}V_{j}||^{2}_{F},$
and we optimize this by alternating between cluster assignments and the JD of each cluster. The algorithm details are in Appendix A.3. Typically, the goal with large $n$ is to have $|C_{j}|$ grow with $n$ as $r$ becomes fixed. Comparing $k$ rank-$r$ $\operatorname{JD-Full}$ clusters to a rank-$kr$ $\operatorname{JD-Full}$ single cluster compression, observe that the clustered approach requires $O(dkr+nr^{2})$ parameters, while the single-cluster approach requires $O(dkr+nk^{2}r^{2})$ parameters due to the increased sizes of the $\Sigma_{i}$s. While these two approaches have the same rank, note that they may have different reconstruction abilities. Empirically, we find that multiple clusters significantly aid performance for $n\geq 100$.
@@ 4 Theoretical Analysis
In this section, we seek to better understand the role of the joint diagonalization method presented in §3.1 and how this understanding further motivates the clustering approach. We will focus on the full-$\Sigma_{i}$ case with orthogonal $U$, $V$ matrices. Note that, for the same $r$, the $r\!\!-\!\!\mathsf{JD\!-\!Diag}$ has at least as large reconstruction error as $r\!\!-\!\!\mathsf{JD\!-\!Full}$ since it imposes an additional constraint on the $\Sigma_{i}$.
Firstly, note that perfect reconstruction can be achieved if and only if $r$ is large enough, since there exist $U,V$ such that all the $B_{i}$, $A_{i}$ are in the spans of $U,V$ resp. if and only if $r\geq\tilde{r}$:
@@
Note $\max_{i}r_{i}\leq\tilde{r}\leq\sum_{i=1}^{n}r_{i}$. Then $\mathsf{JD-Full}$ (equation 2) with $r=\tilde{r}$ achieves lossless compression (perfect reconstruction), and using $r<\tilde{r}$ will give nonzero reconstruction error.
@@
This reconstruction loss can be significant, as the following theorem shows (proved in Appendix B):
@@
Let $\sigma_{j}$ be the singular values of $L$, sorted from largest to smallest, and let $\bar{\sigma}_{j}$ be the singular values of $\sum_{i=1}^{n}B_{i}A_{i}$. Then, using $\mathsf{JD-Full}$ (equation 2),
@@
Remark 2 (Upper bound and grouping).
@@
Suppose (e.g., due to normalization) that the inputs to the joint diagonalization algorithm all have unit Frobenius norm, i.e., $\|B_{i}A_{i}\|_{\mathrm{Fro}}=1$. Moreover, assume that the LoRAs are all orthogonal in the sense $\mathrm{tr}((B_{i}A_{i})(B_{j}A_{j})^{\top})=0$ for $i\neq j$. Then, using the $\mathsf{JD-Full}$ method equation 2, we have $1\leq\sum_{i=1}^{n}\|\Sigma_{i}\|^{2}_{\mathrm{Fro}}\leq\min(r^{2},n),$ implying that the sum of squared Frobenius norms of the reconstructed LoRAs satisfies
@@
Since the loss space of transformers is highly complex, increasing weight reconstruction error does not necessarily imply degrading LLM performance. Interestingly, in Figure 4 below, we see that while large reconstruction error rapidly decreases performance, moderate (but still relatively large, at around 0.4) reconstruction error does not damage performance and may even slightly outperform the zero-error setting. This observation motivates our focus on minimizing weight reconstruction error, while also suggesting that our approach is capable of achieving something deeper than compression. Specifically, the tendency of joint diagonalization is to find subspaces that are shared among many LoRAs when $r$ is large, and to merge subspaces when $r$ is small. When $r$ is particularly small, this tendency towards averaging all or some of the LoRA signals directly connects to the concept of merging LoRAs, whose empirical success (Shah et al., 2023; Huang et al., 2024) could explain the success of our procedure despite the nonlinearity of transformers.
Experiments in Appendix G.9 explore this idea further, comparing reconstruction of real-world LoRAs to reconstruction of randomly sampled LoRAs. The reconstruction error is generally large, but significantly lower than the reconstruction error for random noise, indicating that there is a major shared component between the LoRAs that is being successfully retained.
That said, as the number of LoRAs grows, the shared component may not be significant enough to maintain sufficiently low reconstruction error with low rank $r$. This motivates the introduction of clustering in §3.2, since clustering seeks to find groups of LoRAs that are similar and better compressible by joint diagonalization. In particular, if the number of clusters $|C_{j}|$ grows with $n$, the reconstruction error may no longer degrade with $n$ even when $r$ is fixed.
In the extreme case where $|C_{j}|=n$, each LoRA is compressed independently. By the Eckart-Young Theorem, JD applied to a single LoRA reduces to an SVD, replacing each rank-$r_{i}$ LoRA adapter $B_{i}A_{i}$ with a reduced rank-$r$ approximation, where typically $r<\frac{1}{n}\sum_{i=1}^{n}r_{i}$:
$\operatorname{SVD}_{r}(B_{i}A_{i})=U_{i}\Sigma_{i}V_{i}^{\top},\quad\forall i=1,\dots,n.$ (4)
As $\Sigma_{i}V_{i}^{\top}$ can be saved as a single matrix, this approach has $rn(d_{A}+d_{B})$ parameters. We refer to this $|C_{j}|=n$ method as $\mathsf{r-SVD}$ and find that it underperforms our other methods, while outperforming the baseline uncompressed LoRAs significantly. This result parallels Jiang et al. (2023b)’s observation that lowering LoRA ranks is beneficial for multi-task learning and model merging.